# Crack detection by holomorphic neural networks and transfer-learning-enhanced genetic optimization

### v1

@@ Abstract
A new strategy for detecting cracks in 2D solids based on strain data is introduced. Crack detection is formulated as an inverse problem and solved using genetic optimization. The novelty lies in the evaluation of the model response at each generation. Specifically, the solution to the corresponding plane elasticity problem is expressed via holomorphic potentials, which are determined by training two holomorphic neural networks. As the potentials satisfy equilibrium and traction-free conditions along the crack faces a priori, the training proceeds quickly based solely on boundary information. Training efficiency is further improved by splitting the genetic search into long-range and short-range stages, enabling the use of transfer learning in the latter. The new strategy is tested on three benchmark problems, showing that an optimal number of training epochs exists that provides the best overall performance. A comparison is also made with a popular crack detection approach that uses XFEM to compute the model response. Under the assumption of identical stress-field representation accuracy, the proposed method is found to be between 7 and 23 times faster than the XFEM-based approach. While the strategy is presented here for the simplified case of a single internal crack, generalization is feasible. Overall, the present findings demonstrate that combining genetic optimization with holomorphic neural networks and transfer learning offers a promising avenue for developing crack detection strategies with higher efficiency than those currently available.
@@ 1 Introduction
The extensive amount of data required to train machine learning models has prompted numerous investigations in the rapidly growing field of physics-informed machine learning [15, 16], and in particular, physics-informed neural networks (PINNs) [17, 18]. These approaches require substantially less data, as they rely on enforcing physical constraints. In the simplest form of the PINN approach, a neural network is employed as a global ansatz function to solve a given boundary value problem. The network parameters (weights and biases) are optimized during training by minimizing a loss function that includes the residuals of the governing partial differential equations, the boundary conditions, and any deviations from available experimental data.
@@
Recently, two enrichment strategies have been proposed to further improve the accuracy and efficiency of HNNs for fracture mechanics problems [29]. The first involves enriching the HNNs with the square-root term from Williams’ series, while the second leverages Rice’s global representation of the solution for a straight crack to decouple the holomorphic part of the solution from the singular, non-holomorphic terms. In addition to significantly improving accuracy in the presence of cracks, both strategies were shown to be highly compatible with TL, enabling a ten-fold reduction in training time for repeated computations. Remarkably, preliminary comparisons indicated that the computational efficiency of enriched HNNs when using TL was nearly comparable to that of Abaqus XFEM [29].
In light of the above, enriched HNNs appear to be particularly well suited for crack detection analyses. Accordingly, the present work aims to explore the benefits of combining HNNs and TL with genetic optimization to detect cracks in two-dimensional linear elastic solids based on strain sensor data. The remainder of this article is organized as follows. Section 2 provides the necessary theoretical background, including the complex-valued formulation of plane linear elasticity, the corresponding complex representation of the solution for an internal crack, and an introduction to HNNs. Section 3 defines the crack detection problem and describes two solution strategies: the state-of-the-art approach based on GA-XFEM, and a new approach combining HNNs with GA. A thorough comparison between the two methods is presented through numerical tests in Section 4. Finally, Section 5 summarizes the main findings and conclusions.
@@ 2.3 Solution using holomorphic neural networks
A holomorphic neural network (HNN) is a complex-valued artificial neural network employing holomorphic activation functions. As demonstrated in [27], the choice of holomorphic activation functions – typically the complex exponential – has two important consequences. First, the network output remains holomorphic across the entire complex plane for any values of the network parameters. Second, the network can approximate any holomorphic function to an arbitrary degree of accuracy, provided a sufficient number of layers and neurons is used. These properties make HNNs particularly suitable for solving boundary value problems in which the solution can be expressed in terms of holomorphic potentials [27]. In fact, by leveraging the principles of the PINN approach, HNNs can effectively learn such potentials by minimizing a loss function that quantifies the violation of the boundary conditions. In contrast to the standard PINN approach, no explicit enforcement of the governing differential equations is required, as they are inherently satisfied. This leads to significant gains in both accuracy and computational efficiency [27, 29, 33].
@@ 4 Results
The GA–HNN and the GA–XFEM crack detection strategies were both implemented in Python (link to github rep will be made available upon acceptance of the article). To evaluate the model response, the GA-HNN approach leverages the original PIHNN library [27], while the GA–XFEM approach employs Abaqus XFEM.
@@ 4.1 Convergence analysis
As mentioned in Section˜3.3, the number of training epochs is expected to have a significant impact on the convergence behavior – and thus the efficiency – of the proposed GA–HNN crack detection strategy. To investigate this aspect, the first numerical experiment shown in Figure˜4(a) is considered. The optimization problem is solved using the parameter values listed in Table˜1. The search space is defined as $\mathcal{P}=[-0.8L,+0.8L]\times[-0.8L,+0.8L]$, and a minimum admissible crack length of $0.1L$ is imposed. The model response is evaluated using neural networks consisting of 3 hidden layers with 10 neurons per layer, together with $N_{\sigma}=300$ training points. Given the relative simplicity of the boundary geometry and boundary conditions, this value of $N_{\sigma}$ is sufficient to ensure an accurate computation of the model response, provided that the number of training epochs is large enough.
@@ 4.2 Efficiency comparison
Parameters specific to the GA–HNN strategy are set as described in Section˜4.1. In addition, the hyperparameters controlling the training of the HNNs, like learning rate, optimizer, etc., are set equal to those used in [29]. Regarding XFEM, a regular rectangular mesh and the default Abaqus selection of all the XFEM parameters are adopted. The crack is modeled as a wire feature, and the feature end points are updated each time the model response is to be evaluated, without modifying the underlying mesh.
@@ 5 Conclusions
The present work presents a new strategy for detecting cracks in 2D solids based on strain sensor data. Following previous studies, crack detection is formulated as an inverse problem and solved using genetic optimization. The novelty lies in the evaluation of the model response at each generation. Specifically, the solution to the corresponding plane elasticity problem is expressed in terms of two holomorphic potentials, which are found by training two HNNs. As the potentials automatically satisfy equilibrium and traction-free conditions along the crack faces, the training proceeds quickly based solely on the boundary conditions. Training efficiency is further improved by splitting the genetic algorithm into long-range and short-range stages, enabling the use of transfer learning in the latter.
The new strategy has been assessed through three benchmark tests. The results show that an optimal number of epochs exists for training the HNNs, which provides a balance between accuracy and computational cost and thereby yields the best overall performance. The proposed strategy has also been compared with the reference approach, in which XFEM is used to compute the model response. Using the requirement of identical stress-field representation accuracy to determine both the HNN architecture and the XFEM mesh size, the proposed method is found to be between 7 and 23 times faster than the standard XFEM-based approach. This improvement arises partly from the more efficient two-stage search mechanism, and partly from the more rapid evaluation of the model response enabled by transfer learning.
The present study builds on strong assumptions, as it considers the simplified case of a single internal crack. However, extending the approach to an edge crack would be straightforward, requiring only an update of the expressions for the complex potentials in Equation˜6 [29]. An extension to multiple cracks could also be achieved by leveraging domain decomposition techniques [27].
@@ Acknowledgments
Code will be released publicly following article acceptance.

### v2

@@ Abstract
A physics-informed machine learning framework based on holomorphic neural networks is introduced for detecting cracks in two-dimensional solids from strain or displacement data. Crack detection is formulated as an inverse problem in which the crack size, orientation, and location are treated as unknowns. The problem is solved using genetic optimization, where the fitness function is evaluated by expressing the solution of the corresponding plane elasticity problem in terms of holomorphic potentials, which are then determined through the training of two holomorphic neural networks. As the potentials satisfy equilibrium and traction-free conditions along the crack faces a priori, the training proceeds quickly based solely on boundary information. Training efficiency is further improved by splitting the genetic search into long-range and short-range stages, enabling the use of transfer learning in the latter. The new strategy is tested on three benchmark problems, showing that an optimal number of training epochs exists that provides the best overall performance. A comparison is also made with a popular crack detection approach that uses XFEM to compute the model response. Under the assumption of identical stress-field representation accuracy, the proposed method is found to be between 7 and 23 times faster than the XFEM-based approach. Furthermore, the proposed method appears to be less sensitive to noise in the input data. Overall, the present findings demonstrate that combining genetic optimization with holomorphic neural networks and transfer learning offers a promising avenue for developing crack detection strategies with higher efficiency than those currently available.
@@ 1 Introduction
In an effort to reduce the reliance on experimental training data, several studies have proposed hybrid approaches where machine learning is supported by computational mechanics tools [10]. For instance, Lee et al. [19] proposed a hybrid framework in which deep neural networks are trained on damage scenarios generated from a finite element model updated using experimental modal data. Zhong et al. [41] introduced a phase-space matrix representation of vibration signals combined with convolutional neural networks, achieving high robustness to measurement noise and finite element modeling uncertainty, even with a limited number of sensors. For composite structures, Nishioka et al. [25] developed a defect localization strategy that couples finite element simulated stress fields with graph neural networks, exploiting mesh topology through graph attention mechanisms to predict three-dimensional delamination cracks. In large-scale marine structures, Bardiani et al. [3] integrated finite element modeling with machine learning to enable real-time damage detection and localization under extreme operational conditions, illustrating the scalability of hybrid machine learning–computational mechanics systems for complex geometries.
On a different track, the extensive amount of data required to train machine learning models has also prompted numerous investigations in the rapidly growing field of physics-informed machine learning [15, 32], and in particular, physics-informed neural networks (PINNs) [18, 28]. These approaches require substantially less data, as they rely on enforcing physical constraints. In the simplest form of the PINN approach, a neural network is employed as a global ansatz function to solve a given boundary value problem. The network parameters (weights and biases) are optimized during training by minimizing a loss function that includes the residuals of the governing partial differential equations, the boundary conditions, and any deviations from available experimental data.
@@
Recently, two enrichment strategies have been proposed to further improve the accuracy and efficiency of HNNs for fracture mechanics problems [6]. The first involves enriching the HNNs with the square-root term from Williams’ series, while the second leverages Rice’s global representation of the solution for a straight crack to decouple the holomorphic part of the solution from the singular, non-holomorphic terms. In addition to significantly improving accuracy in the presence of cracks, both strategies were shown to be highly compatible with TL, enabling a ten-fold reduction in training time for repeated computations. Remarkably, preliminary comparisons indicated that, for comparable levels of target accuracy, the computational efficiency of enriched HNNs when using TL was nearly comparable to that of Abaqus XFEM [6].
In light of the above, the present work proposes a novel and efficient framework for crack detection in two-dimensional linear elastic solids that leverages HNNs and transfer learning. This framework combines the enriched HNN formulation introduced by Calafà et al. [6] with a modified genetic optimization algorithm in which the crossover and mutation search stages are decoupled. The remainder of this article is organized as follows. Section 2 provides the necessary theoretical background, including the complex-valued formulation of plane linear elasticity, the corresponding complex representation of the solution for an internal crack, and an introduction to HNNs. Section 3 defines the crack detection problem and describes two solution strategies: the state-of-the-art approach based on GA-XFEM, and the new approach combining HNNs with GA. The performance of the new approach is thoroughly analyzed in Section 4 in terms of convergence, noise sensitivity, and efficiency in relation to the standard GA-XFEM approach. Finally, Section 5 summarizes the main findings and conclusions.
@@ 2.3 Solution using holomorphic neural networks
A holomorphic neural network (HNN) is a complex-valued artificial neural network employing holomorphic activation functions. As demonstrated in [5], the choice of holomorphic activation functions – typically the complex exponential – has two important consequences. First, the network output remains holomorphic across the entire complex plane for any values of the network parameters. Second, the network can approximate any holomorphic function to an arbitrary degree of accuracy, provided a sufficient number of layers and neurons is used. These properties make HNNs particularly suitable for solving boundary value problems in which the solution can be expressed in terms of holomorphic potentials [5]. In fact, by leveraging the principles of the PINN approach, HNNs can effectively learn such potentials by minimizing a loss function that quantifies the violation of the boundary conditions. In contrast to the standard PINN approach, no explicit enforcement of the governing differential equations is required, as they are inherently satisfied. This leads to significant gains in both accuracy and computational efficiency [5, 6, calafà2026].
@@ 4 Results
The GA–HNN and the GA–XFEM crack detection strategies were both implemented in Python [^1] https://github.com/jonashund/holomorphic_crack_detection. To evaluate the model response, the GA-HNN approach leverages the original PIHNN library [5], while the GA–XFEM approach employs Abaqus XFEM.
@@ 4.1 Convergence analysis
As mentioned in Section˜3.3, the number of training epochs is expected to have a significant impact on the convergence behavior – and thus the efficiency – of the proposed GA–HNN crack detection strategy. To investigate this aspect, the first numerical experiment shown in Figure˜4(a) is considered. The optimization problem is solved using the parameter values listed in Table˜1. The search space is defined as $\mathcal{P}=[-0.8L,+0.8L]\times[-0.8L,+0.8L]$, and a minimum admissible crack length of $0.1L$ is imposed. The model response is evaluated by training two holomorphic neural networks, each composed of three fully connected hidden layers with ten neurons per layer. The values of the main hyperparameters used for the training are reported in Table˜1. A total of $N_{\sigma}=300$ training points are employed, which are uniformly distributed along the plate boundary. Given the relative simplicity of the boundary geometry and boundary conditions, this value of $N_{\sigma}$ is sufficient to ensure an accurate computation of the model response, provided that the number of training epochs is large enough.
@@
Optimizer Learning rate $s_{u}$ $s_{\sigma}$ $N_{\sigma}$
ADAM $0^{-2} 11$ 11 300300
Table 2: Main hyperparameters used to train the holomorphic neural networks.
@@ 4.2 Efficiency comparison
Parameters specific to the GA–HNN strategy, including the hyperparameters controlling the training of the HNNs, are set as described in Section˜4.1. Regarding XFEM, a regular rectangular mesh and the default Abaqus selection of all the XFEM parameters are adopted. The crack is modeled as a wire feature, and the feature end points are updated each time the model response is to be evaluated, without modifying the underlying mesh.
@@
4.3 Noise sensitivity
The results presented in the previous sections are based on the assumption that the strain values measured by the sensors are error free. However, in real-world scenarios, strain data are inevitably corrupted by noise to some extent. It is therefore important to analyze how the proposed GA–HNN crack detection strategy is affected by the presence of noise in the strain measurements.
[…]
In the above expression, the first term on the right-hand side represents the reference, noise-free strain. The second term corresponds to the root-mean-square value of the noise-free strain taken over all sensors, multiplied by a normally distributed random variable $Y$ with zero mean and unit variance, and by a real-valued parameter $m$ that controls the noise level. Both the GA–HNN and GA–XFEM algorithms are run on the noisy strain data for a fixed number of 50 generations. The resulting level of accuracy is then quantified by computing the crack detection error, defined as:
$E=|z^{+}_{T}-z^{+}_{F}|+|z^{-}_{T}-z^{-}_{F}|$ (20)
where $z^{+}_{T},z^{-}_{T}$ are the locations of the right and left tips of the true crack, and $z^{+}_{F},z^{-}_{F}$ are the corresponding tip locations for the crack in the population that exhibits the best fitness. The procedure is repeated for values of $m$ equal to 0%, 5%, 10% and 15%, and for each $m$ value, 10 independent runs of the algorithms are considered to get statistical insight.
Figure˜10 presents the evolution of the crack detection error as a function of the noise level. For both the GA–HNN and GA–XFEM strategies, the data are normalized by the error obtained in the noise-free case, in order to highlight the relative effect of noise. As expected, the crack detection error increases monotonically with increasing noise level, since higher noise renders the inverse problem less well posed. For a noise level of 15%, the error is approximately 210% larger than in the noise-free case for the GA–HNN approach, and about 270% larger for the GA–XFEM approach. In this respect, it is worth noting that the GA–HNN strategy is less affected by noise for all noise levels considered. This consistent behavior suggests that the GA–HNN approach may be intrinsically more robust to noise than GA–XFEM. This improved robustness may be attributed to the different manner in which the model response is evaluated, with GA–HNN relying on an approximation of the response, as discussed in Section˜4.1.
Figure 10: Impact of strain measurement noise on the accuracy of the GA–HNN and GA–XFEM crack detection strategies. For each noise level, the accuracy is expressed in terms of the crack detection error defined in Equation˜20. The reported values correspond to the mean over ten independent runs and are normalized by the mean error obtained in the noise-free case.
@@ 5 Conclusions
A physics-informed machine learning framework based on holomorphic neural networks is introduced for detecting cracks in two-dimensional solids from strain or displacement data. Crack detection is formulated as an inverse problem and solved using genetic optimization. At each generation, the fitness function is evaluated by expressing the solution to the corresponding elasticity problem in terms of two holomorphic potentials, which are subsequently found by training two HNNs. As the potentials automatically satisfy equilibrium and traction-free conditions along the crack faces, the training proceeds quickly based solely on the boundary conditions. Training efficiency is further improved by splitting the genetic algorithm into long-range and short-range stages, enabling the use of transfer learning in the latter.
The new strategy has been assessed through three benchmark tests. The results show that an optimal number of epochs exists for training the HNNs, which provides a balance between accuracy and computational cost and thereby yields the best overall performance. The proposed strategy has also been compared with the reference approach, in which XFEM is used to compute the model response. Using the requirement of identical stress-field representation accuracy to determine both the HNN architecture and the XFEM mesh size, the proposed method is found to be between 7 and 23 times faster than the standard XFEM-based approach. This improvement arises partly from the more efficient two-stage search mechanism, and partly from the more rapid evaluation of the model response enabled by transfer learning. Furthermore, the proposed strategy appears to exhibit greater robustness to strain measurement noise than the standard XFEM-based approach.
The present study builds on strong assumptions, as it considers the simplified case of a single internal crack. Extending the approach to an edge crack would be straightforward, requiring only an update of the expressions for the complex potentials in Equation˜6 [6]. Extension to multiple cracks would be more challenging instead. While domain decomposition techniques [5] can, in principle, enable the HNN approach to compute the elastic response of a body containing multiple cracks, such a decomposition requires a priori knowledge of the total number and approximate locations of the cracks, which are typically unknown in multi-crack detection problems. Nevertheless, approaches based on multiscale algorithms [30] and the introduction of topological variables [29] may provide viable strategies to overcome this limitation.
@@ Acknowledgments
The Python code associated with the present work is available at https://github.com/jonashund/holomorphic_crack_detection.